# Thermodynamic Limit in Learning Period Three

### v1

@@ Abstract
A continuous one-dimensional map with period three includes all periods. This raises the following question: Can we obtain any types of periodic orbits solely by learning three data points? We consider learning period three with random neural networks and report the universal property associated with it. We first show that the trained networks have a thermodynamic limit that depends on the choice of target data and network settings. Our analysis reveals that almost all learned periods are unstable and each network has its characteristic attractors (which can even be untrained ones). Here, we propose the concept of characteristic bifurcation expressing embeddable attractors intrinsic to the network, in which the target data points and the scale of the network weights function as bifurcation parameters. In conclusion, learning period three generates various attractors through characteristic bifurcation due to the stability change in latently existing numerous unstable periods of the system.
With the advent of reservoir computing (RC) [1], which exploits the dynamics of high-dimensional dynamical systems for learning, many interesting properties of random networks have been discovered from a dynamical systems perspective. For example, with a fixed-weight random recurrent neural network, the so-called echo-state network (ESN) [2]—as a reservoir—we can create an autonomous system that emulates target chaotic systems by simply fitting its readout layer [3, 4]. Such an attractor-embedding ability is linked to the existence of the generalized synchronization between an input dynamical system and a reservoir [4, 5, 6, 7]. In addition, recent studies have revealed that RC can simultaneously embed multiple attractors [8] and may have untrained attractors that are not part of the training data [8, 9, 10, 11, 12]. Surprisingly, a successful reservoir computer only needs a few pairs of bifurcation parameter values and corresponding trajectories to reconstruct the entire bifurcation structure of a target system [10, 11]. These properties are valuable, for example, in the context of robot locomotion control using dynamical system attractors that can significantly reduce the training data [13, 14, 15]. Thus, the powerful generalization and multifunctionality aspects of RC are related to the dynamical systems properties of a learning machine.
@@
We write $m\succ n$ whenever $m$ is to the left of $n$ in Eq. (1). As a consequence of both theorems, a continuous one-dimensional map with period three has all periods. We note that such a map also has an uncountable set containing no (asymptotically) periodic points, the so-called scrambled set, and is chaotic in the sense of Li–Yorke [18].
Figure 1: Trained maps $f_{N}^{*}$ for activation functions $\phi=\text{erf}$ (top left), $\sin$ (top right), $\cos$ (bottom left), and ReLU (bottom right), with target period three ${\cal D}=\{-1,1,0\}$ and input intensity $\sigma=1.0$. The blue-colored areas indicate the maximum–minimum regions of $f_{N}^{*}$ for 100 different realizations. The red circles and the red dotted lines show ${\cal D}$ and the thermodynamic limit $f_{\infty}^{*}$, respectively. The darkness of blue corresponds to the number of nodes $N$, describing that $f_{N}^{*}$ degenerates into $f_{\infty}^{*}$ as $N$ increases.
@@
Now, the following natural question arises: Can we obtain all the periods in the network through training only period three? If we successfully train period three in one-dimensional dynamics, then the straightforward answer is “yes.” However, this question is somewhat naive, since the above theorems do not reveal whether the obtained periods are stable. Instead, we should ask the following question: Which kind of stable orbits (attractors) can we obtain by learning period three? This paper is devoted to theoretically answer this question in terms of all aspects. Here, we consider training readout weights of a two-layer random feedforward neural network to learn period three [19, 20, 21, 22]: $f_{N}(x)\equiv\frac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{i}^{\text{out}}\phi\left(h_{i}(x)\right)$, $h_{i}(x)\equiv W_{i}^{\text{in}}x+b_{i}^{\text{in}}$, where $x\in\mathbb{R}$ is an input of the network, $W^{\text{in}}\in\mathbb{R}^{N\times 1}$ and $b^{\text{in}}\in\mathbb{R}^{N}$ are the input weights and biases, respectively, randomly drawn from a Gaussian distribution with zero mean and variance given by $\sigma^{2}$, $\phi\mathbin{:}\mathbb{R}\rightarrow\mathbb{R}$ is an activation function, and $W^{\text{out}}\in\mathbb{R}^{1\times N}$ is the output weights matrix optimized by a learning method described subsequently (see Sec. LABEL:S:sec:TL in the Supplementary Material [23] for the other network settings and detailed derivations, which includes Refs. [24, 25, 26]). We denote the trained network output by $f_{N}^{*}(x)$ and call $\sigma$ the input intensity. Hereafter, we consider the following specific activation functions: bounded and smooth $\phi=\text{erf},\sin,\cos$ and unbounded and non-smooth $\phi=\text{ReLU}$. Actually, this model can be regarded as a special case of an ESN in which the spectral radius of the adjacency matrix is zero. With the trained network $f_{N}^{*}$, we study the dynamical system $x_{n+1}=f_{N}^{*}(x_{n})$ created from a closed loop that connects its input and output. The attractors of $f_{N}^{*}$ should depend on the network structure, including the realizations of $W^{\text{in}}$ and $b^{\text{in}}$ and the choice of $\phi$, which makes the above question non-trivial.
@@
Our findings are twofold. First, concerning general training data, we show that the trained map $f_{N}^{*}$ degenerates into its thermodynamic limit $f_{\infty}^{*}$ as $N$ increases (Fig. 1) based on the neural tangent kernel (NTK) theory [27, 28] under certain assumptions. Second, in learning period three, we reveal that in contrast to our expectation, almost all acquired periods are unstable; it has characteristic attractors corresponding to the choice of target period three, nonlinearity $\phi$, and input intensity $\sigma$. Interestingly, the changes in the target data points and $\sigma$, which can be regarded as controllable system parameters, cause the bifurcation of the characteristic attractors, enabling learning machines to embed a diverse range of dynamics in their bifurcation structure. We discuss the range of applicability of our proposed approach toward physical learning machines at the end of this paper.
@@
We denote the training dataset and its number by ${\cal D}\subseteq\mathbb{R}\times\mathbb{R}$ and $|{\cal D}|$, respectively, and assume $N\geq|{\cal D}|$. We use ${\cal X}$ and ${\cal Y}$ vectors to denote the input and output data and define them as ${\cal X}\equiv[x_{1},\ldots,x_{|{\cal D}|}],{\cal Y}\equiv[y_{1},\ldots,y_{|{\cal D}|}]\in\mathbb{R}^{|{\cal D}|}$, where $(x_{i},y_{i})\in{\cal D}$. In learning period three ($a\mapsto b\mapsto c\mapsto a\mapsto\cdots$), the target input–output pairs are ${\cal D}=\left\{(a,b),(b,c),(c,a)\right\}$, ${\cal X}=[a,b,c]$, and ${\cal Y}=[b,c,a]$. For the sake of simplicity, we write ${\cal D}=\{a,b,c\}$ and assume $a<b$. Note that a period-3 orbit is of two types: $\{a,b,c\}$ and $\{b,a,c\}$. Generally, a period-$n$ orbit has $(n-1)!$ types. For a given ${\cal D}$, we optimize $W^{\text{out}}$ by minimizing a squared error via continuous-time gradient descent (GD) to easily take the limit $N\rightarrow\infty$ [27, 28, 29, 30].
We denote by $\theta_{t}\in\mathbb{R}^{N}$ and $f_{t}(x,\theta)\in\mathbb{R}$ the time-dependence of the vectorized output weights and $f_{N}(x)$ at time $t$, respectively. The dynamics of GD with the initialization $\theta_{0}=\bm{0}$ yield the following formulae of the optimal output weights $\left(W^{\text{out}}\right)^{*}$ and $f_{N}^{*}(x)$:
$\theta_{t}=\nabla_{\theta}f_{0}({\cal X})^{\top}\hat{\Theta}_{0}^{-1}\left(I-e^{-\eta\hat{\Theta}_{0}t}\right){\cal Y},$
$\left(W^{\text{out}}\right)^{*\top}=\lim_{t\rightarrow\infty}\theta_{t}=\nabla_{\theta}f_{0}({\cal X})^{\top}\hat{\Theta}_{0}^{-1}{\cal Y},$ (2)
$f_{t}(x,\theta)=\hat{\Theta}_{0}(x,{\cal X})\hat{\Theta}_{0}^{-1}\left(I-e^{-\eta\hat{\Theta}_{0}t}\right){\cal Y},$
$f_{N}^{*}(x)=\lim_{t\rightarrow\infty}f_{t}(x,\theta)=\hat{\Theta}_{0}(x,{\cal X})\hat{\Theta}_{0}^{-1}{\cal Y},$ (3)
@@
where $\eta>0$ is a learning rate and $\nabla_{\theta}f_{0}({\cal X})\in\mathbb{R}^{|{\cal D}|\times N}$, $\hat{\Theta}_{0}(x,{\cal X})\in\mathbb{R}^{1\times|{\cal D}|}$, and $\hat{\Theta}_{0}\in\mathbb{R}^{|{\cal D}|\times|{\cal D}|}$ are the matrices; these matrices are given in the following manner:
$\left.\nabla_{\theta}f_{0}({\cal X})\right._{ij}=\frac{1}{\sqrt{N}}\phi\left(h_{j}(x_{i})\right),\,\left.\hat{\Theta}_{0}(x,{\cal X})\right._{i}\equiv\hat{\Theta}_{0}\left(x,x_{i}\right),$
$\left.\hat{\Theta}_{0}\right.\equiv\left.\hat{\Theta}_{0}({\cal X},{\cal X})\right.\equiv\nabla_{\theta}f_{0}({\cal X})\nabla_{\theta}f_{0}({\cal X})^{\top},$ (4)
$\hat{\Theta}_{0}(x,y)=\frac{1}{N}\sum_{i=1}^{N}\phi\left(h_{i}(x)\right)\phi\left(h_{i}(y)\right).$ (5)
We note that $\hat{\Theta}_{0}$ is symmetric and assumed to be positive definite. Eqs. (2) and (4) yield
$\left(W^{\text{out}}\right)^{*\top}=\nabla_{\theta}f_{0}({\cal X})^{\top}\left(\nabla_{\theta}f_{0}({\cal X})\nabla_{\theta}f_{0}({\cal X})^{\top}\right)^{-1}{\cal Y},$
implying that $\theta_{0}=\bm{0}$ will lead $W^{\text{out}}$ to the minimum norm solution (MNS) [29, 31, 30, 32], which is given by the pseudoinverse of $\nabla_{\theta}f_{0}({\cal X})$ (“extreme learning machine” [33, 20, 21]).
@@
By the law of large numbers, $\hat{\Theta}_{0}(x,y)$ (Eq. (5)) converges in probability to $\Theta(x,y)$—that is the expectation over random variables $[\omega,\beta]\sim{\cal N}(\bm{0},\sigma^{2}I)$—in the limit $N\rightarrow\infty$ because the components of $W^{\text{in}}$ and $b^{\text{in}}$ follow an iid Gaussian distribution [27]:
@@
$\Theta(x,y)$ is called the NTK [27]. Defining $\Theta(x,{\cal X})$ and $\Theta$ in the same manner as $\hat{\Theta}_{0}(x,y)$, we acquire $f^{*}_{\infty}(x)$, since $f_{N}^{*}(x)$ (Eq. (3)) is calculated from only the values of $\hat{\Theta}_{0}(x,y)$:
@@
where we again assume $\Theta$ to be positive definite. Moreover, the output of infinite-width ESN, which is described by recurrent kernels [34, 35, 36], reduces to the form of Eq. (7) in the limit where the regularization parameter goes to zero without considering the bias term of the network output.
With regard to learning period three, $f_{\infty}^{*}$ is given by
@@
To investigate the dynamics of $f_{\infty}^{*}$, we compute the trajectory $\left\{x_{n}\right\}_{n=0}^{T}$ of $T+1\gg 1$ steps with an initial state $x_{0}$ and calculate the Lyapunov exponent and period of attractors from trajectories. The Lyapunov exponent expresses the sensitivity of a dynamical system to initial conditions and is calculated using the following equation:
$\lambda_{T}=\frac{1}{T}\sum_{n=1}^{T}\ln{\left|\frac{d}{dx}f_{\infty}^{*}(x_{n})\right|}.$ (9)
@@
$\frac{d}{dx}f_{\infty}^{*}(x)=\frac{\partial}{\partial x}\Theta(x,{\cal X})\Theta^{-1}{\cal Y}.$ (10)
@@
We also calculate the period of attractors as the minimum integer $n\in[1,n_{\text{max}}]$ that satisfies the following inequality ($n_{\text{max}}=10\text{ or }20$):
@@
$&\frac{\left|(f_{\infty}^{*})^{n}(x_{T})-x_{T}\right|}{\max\left\{\left|x_{T}\right|,\left|(f_{\infty}^{*})^{n}(x_{T})\right|\right\}}\leq 10^{-12},\\ \text{where}\quad&(f_{\infty}^{*})^{1}\equiv f_{\infty}^{*},\,(f_{\infty}^{*})^{k+1}\equiv f_{\infty}^{*}\circ(f_{\infty}^{*})^{k}.\end{aligned}\right.$ (11)
@@
For $\phi=\text{erf},\sin,\cos,\text{ReLU}$, there exist the analytical solutions of $\Theta(x,y)$ (Sec. LABEL:S:subsec:NTK in the Supplementary Material [23], which includes Refs. [37, 38]):
$&\Theta^{\text{erf}}(x,y)\\ &=\frac{2}{\pi}\arcsin\frac{2\sigma^{2}(1+xy)}{\sqrt{\left[1+2\sigma^{2}(1+x^{2})\right]\left[1+2\sigma^{2}(1+y^{2})\right]}},\end{aligned}$ (12)
$\Theta^{\sin}(x,y)=\frac{1}{2}\left\{e^{-\frac{\sigma^{2}}{2}(x-y)^{2}}-e^{-\frac{\sigma^{2}}{2}(x+y)^{2}-2\sigma^{2}}\right\},$ (13)
$\Theta^{\cos}(x,y)=\frac{1}{2}\left\{e^{-\frac{\sigma^{2}}{2}(x-y)^{2}}+e^{-\frac{\sigma^{2}}{2}(x+y)^{2}-2\sigma^{2}}\right\},$ (14)
$\Theta^{\text{relu}}(x,y)=&\frac{\sigma^{2}}{2\pi}\sqrt{\left(1+x^{2}\right)\left(1+y^{2}\right)}\\ &\left\{\sqrt{1-\cos^{2}\psi}+(\pi-\psi)\cos\psi\right\},\end{aligned}$ (15)
@@
where $\psi\equiv\arccos\frac{1+xy}{\sqrt{\left(1+x^{2}\right)\left(1+y^{2}\right)}}$. We note that the trajectories for a bounded activation function—such as $\phi=\text{erf},\sin,\cos$—are also bounded since $\Theta(x,y)$ (Eq. (6)) is the expectation of the product of $\phi$. Moreover, as $\phi=\text{erf},\sin,\cos$ and their corresponding NTKs are analytic, $f_{N}^{*}$ and $f_{\infty}^{*}$ have finitely many attractors, according to Theorem B in Ref. [39] (Sec. LABEL:S:sec:finiteness in the Supplementary Material [23], which includes Ref. [40]). In contrast, certain trajectories for $\phi=\text{ReLU}$, whose NTK is unbounded, will head toward infinity.
@@
Figure 2: Change in dynamical system $f_{\infty}^{*}$ for $\phi=\text{erf}$ with respect to $c$, with $a=-1$, $b=1$, and $\sigma=1.0$. The trajectory $\{x_{n}\}_{n=0}^{T}$ of $T=10^{5}$ steps is computed with different $x_{0}$ and $c$ in given intervals, excluding $c=a,b$. (a) Bifurcation diagram of the characteristic attractors calculated with $-10\leq x_{0}\leq 10$. The dotted lines indicate ${\cal D}$; the diagonal line corresponds to a varying $c$. (b) Change in the basin of attraction where $f_{\infty}^{*}$ has multiple untrained attractors. (c) Change in the map $f_{\infty}^{*}$ in $-0.8\leq c\leq 0.8$. The circles indicate ${\cal D}$ and the red circle indicates a c-independent point $(a,b)=(-1,1)$. As $c$ approaches $a$ or $b$, the folding of $f_{\infty}^{*}$ around ${\cal D}$ becomes larger, making the characteristic attractors wider.

### v2

@@ Abstract
A continuous one-dimensional map with period three includes all periods. This raises the following question: Can we obtain any types of periodic orbits solely by learning three data points? In this letter, we report the answer to be yes. Considering a random neural network in its thermodynamic limit, we show that under certain conditions, learning period three can embed attractors with all periods into the network as a bifurcation after learning. The associated universality is explained by a topological conjugacy between the trained network and the classical logistic map.
With the advent of reservoir computing (RC) [1], which exploits the dynamics of high-dimensional dynamical systems for learning, many interesting properties of random networks have been discovered from a dynamical systems perspective. For example, with a fixed-weight random recurrent neural network, the so-called echo-state network (ESN) [2]—as a reservoir—we can create an autonomous system that emulates target chaotic systems by simply fitting its readout layer [3, 4]. Such an attractor-embedding ability is linked to the existence of the generalized synchronization between an input dynamical system and a reservoir [4, 5, 6, 7]. In addition, recent studies have revealed that RC can simultaneously embed multiple attractors [8] and may have untrained attractors that are not part of the training data [8, 9, 10, 11, 12, 13, 14, 15]. Surprisingly, a successful reservoir computer only needs a few pairs of bifurcation parameter values and corresponding trajectories to reconstruct the entire bifurcation structure of a target system [10, 11, 12, 13, 14]. These properties are valuable, for example, in the context of robot locomotion control using dynamical system attractors that can significantly reduce the training data [16, 17, 18]. Thus, the powerful generalization and multifunctionality aspects of RC are related to the dynamical systems properties of a learning machine.
@@
We write $m\succ n$ whenever $m$ is to the left of $n$ in Eq. (1). As a consequence of both theorems, a continuous one-dimensional map with period three has all periods.
Now, the following natural question arises: Can we obtain all the periods in the network through training only period three? If we successfully train period three in one-dimensional dynamics, then the straightforward answer is “yes.” However, this question is somewhat naive, since the above theorems do not reveal whether the obtained periods are stable. Instead, we should ask the following question: Which kind of stable orbits (attractors) can we obtain by learning period three (LP3)? This paper is devoted to theoretically answer this question in terms of all aspects.
@@
To clarify what LP3 implies, we need to create an one-dimensional recurrent neural network with target period three. LP3 with ESN, as in the standard RC scheme, will violate the assumption of the above two theorems due to the high-dimensionality of the resulting dynamics. Thus, here, we consider training the readout weights of a two-layer random feedforward neural network to learn period three [22, 23, 24, 25]: $f_{N}(x)\equiv\frac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{i}^{\text{out}}\phi\left(h_{i}(x)\right)$, $h_{i}(x)\equiv W_{i}^{\text{in}}x+b_{i}^{\text{in}}$, where $x\in\mathbb{R}$ is an input of the network; $W^{\text{in}}\in\mathbb{R}^{N\times 1}$ and $b^{\text{in}}\in\mathbb{R}^{N}$ are the input weights and biases, respectively, randomly drawn from a Gaussian distribution with zero mean and variance given by $\sigma^{2}$; $\phi\mathbin{:}\mathbb{R}\rightarrow\mathbb{R}$ is an activation function; and $W^{\text{out}}\in\mathbb{R}^{1\times N}$ is the output weights matrix optimized by a learning method described subsequently (see Sec. LABEL:S:subsec:derivation and Sec. LABEL:S:subsec:uniform in the Supplementary Material (SM) [26] for networks with output bias and uniform distribution). Actually, this model can be regarded as a special case of an ESN in the limit where the spectral radius of the adjacency matrix goes to zero. In addition, we denote the trained network output by $f_{N}^{*}(x)$ and call $\sigma$ the scale of weights. With the trained network $f_{N}^{*}$, we study the dynamical system $x_{n+1}=f_{N}^{*}(x_{n})$ created from a closed loop that connects its input and output. The attractors of $f_{N}^{*}$ should depend on the network structure, including the realizations of $W^{\text{in}}$ and $b^{\text{in}}$ and the choice of $\phi$, which makes the above question non-trivial. Hereafter, we consider the following specific activation functions: bounded and smooth $\phi=\text{erf},\sin,\cos$, and unbounded and non-smooth $\phi=\text{ReLU}$.
Figure 1: Trained maps $f_{N}^{*}$ for activation functions $\phi=\text{erf}$ (top left), $\sin$ (top right), $\cos$ (bottom left), and ReLU (bottom right), with target period three ${\cal D}=\{-1,1,0\}$ and scale of weights $\sigma=1.0$. The blue-colored areas indicate the maximum–minimum regions of $f_{N}^{*}$ for 100 different realizations. The red circles and the red dotted lines show ${\cal D}$ and the thermodynamic limit $f_{\infty}^{*}$, respectively. The darkness of blue corresponds to the number of nodes $N$, describing that $f_{N}^{*}$ degenerates into $f_{\infty}^{*}$ as $N$ increases.
@@
Our findings are threefold. First, concerning general training data, we show that the trained map $f_{N}^{*}$ degenerates into its thermodynamic limit $f_{\infty}^{*}$ as $N$ increases (Fig. 1) under certain assumptions. Second, in LP3, we reveal that almost all acquired periods are unstable; it has characteristic attractors corresponding to the choice of target period three, nonlinearity $\phi$, and scale of weights $\sigma$. Third, we prove that neural networks with certain activations can externalize all the latently acquired unstable periods through LP3 as attractors after learning, as a universal bifurcation with regard to the feedback strength. We discuss the range of applicability of our proposed approach to physical learning machines at the end of this paper.
@@
We denote the training dataset and its number by ${\cal D}\subseteq\mathbb{R}\times\mathbb{R}$ and $|{\cal D}|$, respectively, and assume $N\geq|{\cal D}|$. We use ${\cal X}$ and ${\cal Y}$ vectors to denote the input and output data and define them as ${\cal X}\equiv[x_{1},\ldots,x_{|{\cal D}|}],{\cal Y}\equiv[y_{1},\ldots,y_{|{\cal D}|}]\in\mathbb{R}^{|{\cal D}|}$, where $(x_{i},y_{i})\in{\cal D}$. In LP3 ($a\mapsto b\mapsto c\mapsto a\mapsto\cdots$), the target input–output pairs are ${\cal D}=\left\{(a,b),(b,c),(c,a)\right\}$, ${\cal X}=[a,b,c]$, and ${\cal Y}=[b,c,a]$. For the sake of simplicity, we write ${\cal D}=\{a,b,c\}$ and assume $a<b$. Note that a period-3 orbit is of two types: $\{a,b,c\}$ and $\{b,a,c\}$. Generally, a period-$n$ orbit has $(n-1)!$ types. For a given ${\cal D}$, we optimize $W^{\text{out}}$ by least square regression with a minimum norm solution (“extreme learning machine” [27, 23, 24]) or “ridgeless” regression [28, 29, 30, 31]:
$f_{N}^{*}(x)=\lim_{\lambda\searrow 0}\hat{\Theta}(x,{\cal X})\left(\hat{\Theta}+\lambda I_{|{\cal D}|}\right)^{-1}{\cal Y},$ (2)
@@
where $\hat{\Theta}(x,{\cal X})\in\mathbb{R}^{1\times|{\cal D}|}$ and $\hat{\Theta}\equiv\hat{\Theta}({\cal X},{\cal X})\in\mathbb{R}^{|{\cal D}|\times|{\cal D}|}$ are the matrices given in the following manner:
$\left.\hat{\Theta}(x,{\cal X})\right._{i}\equiv\hat{\Theta}\left(x,x_{i}\right),\left.\hat{\Theta}\right._{ij}\equiv\hat{\Theta}\left(x_{i},x_{j}\right),$ (3)
$\hat{\Theta}(x,y)\equiv{\cal R}(x)^{\top}{\cal R}(y)=\frac{1}{N}\sum_{i=1}^{N}\phi\left(h_{i}(x)\right)\phi\left(h_{i}(y)\right),$ (4)
${\cal R}(x)_{i}\equiv\frac{1}{\sqrt{N}}\phi\left(h_{i}(x)\right),\quad{\cal R}(x)\in\mathbb{R}^{N}.$ (5)
We note that Eq. (2) is equivalent to fit $W^{\text{out}}$ with the pseudoinverse of the matrix of hidden states ${\cal R}({\cal X})\equiv[{\cal R}(x_{1})\cdots{\cal R}(x_{|{\cal D}|})]\in\mathbb{R}^{N\times|{\cal D}|}$ [27]: $\left(W^{\text{out}}\right)^{*}={\cal Y}^{\top}{\cal R}({\cal X})^{+}$. If the matrix $\hat{\Theta}$ has full rank, Eq. (2) has the following closed-form expression:
$f_{N}^{*}(x)=\hat{\Theta}(x,{\cal X})\hat{\Theta}^{-1}{\cal Y}.$ (6)
@@
By the law of large numbers, $\hat{\Theta}(x,y)$ (Eq. (4)) converges in probability to $\Theta(x,y)$—that is the expectation over random variables $[\omega,\beta]\sim{\cal N}(\bm{0},\sigma^{2}I)$—in the limit $N\rightarrow\infty$ because the components of $W^{\text{in}}$ and $b^{\text{in}}$ follow an iid Gaussian distribution [32]:
@@
In our model, $\Theta(x,y)$ coincides with the neural tangent kernel (NTK) and the neural network gaussian process (NNGP) kernel [32, 33]. Defining $\Theta(x,{\cal X})$ and $\Theta$ in the same manner as $\hat{\Theta}(x,y)$, we acquire $f^{*}_{\infty}(x)$, since $f_{N}^{*}(x)$ (Eq. (6)) is calculated from only the values of $\hat{\Theta}(x,y)$:
@@
where we again assume $\Theta$ to have full rank. This assumption is valid if $\phi$ is a non-polynomial continuous function and ${\cal X}$ consists of $|{\cal D}|$ distinct points, according to Theorem 3 in Ref. [34] (Sec. LABEL:S:subsec:gram in the SM [26], which includes Ref. [35]). In addition, the output of the infinite-width ESN also reduces to a form similar to that of Eq. (8) with time-varying recurrent kernels [36, 37, 38].
With regard to LP3, $f_{\infty}^{*}$ is given by
@@
To investigate the dynamics of $f_{\infty}^{*}$, we compute the trajectory $\left\{x_{n}\right\}_{n=0}^{T}$ of $T+1\gg 1$ steps with an initial state $x_{0}$ and calculate the Lyapunov exponent and period of attractors from the trajectories. The Lyapunov exponent expresses the sensitivity of a dynamical system to initial conditions and is calculated using the following equation:
$\lambda_{T}=\frac{1}{T}\sum_{n=1}^{T}\ln{\left|\frac{df_{\infty}^{*}}{dx}(x_{n})\right|}.$ (10)
@@
$\frac{df_{\infty}^{*}}{dx}(x)=\frac{\partial\Theta}{\partial x}(x,{\cal X})\Theta^{-1}{\cal Y}.$ (11)
@@
We also calculate the period of attractors as the minimum integer $n\in[1,n_{\text{max}}]$ that satisfies the following inequality ($n_{\text{max}}=10\text{ or }20$ and $\varepsilon=10^{-12}$ except as otherwise noted):
@@
$&\frac{\left|(f_{\infty}^{*})^{n}(x_{T})-x_{T}\right|}{\max\left\{\left|x_{T}\right|,\left|(f_{\infty}^{*})^{n}(x_{T})\right|\right\}}\leq\varepsilon,\\ \text{where}\quad&(f_{\infty}^{*})^{1}\equiv f_{\infty}^{*},\,(f_{\infty}^{*})^{k+1}\equiv f_{\infty}^{*}\circ(f_{\infty}^{*})^{k}.\end{aligned}\right.$ (12)
@@
For $\phi=\text{erf},\sin,\cos,\text{ReLU}$, there exist the analytic solutions of $\Theta(x,y)$. Following Williams [39], the NTK for $\phi=\text{erf}$ is expressed as follows (see Sec. LABEL:S:subsec:NTK in the SM [26] regarding the NTKs for $\phi=\sin,\cos,\text{ReLU}$, which includes Ref. [40]):
$\Theta^{\text{erf}}(x,y)$ (13)
$=\frac{2}{\pi}\arcsin\frac{2\sigma^{2}(1+xy)}{\sqrt{\left[1+2\sigma^{2}(1+x^{2})\right]\left[1+2\sigma^{2}(1+y^{2})\right]}}.$
@@
We note that the trajectories for a bounded activation function—such as $\phi=\text{erf},\sin,\cos$—are also bounded since $\Theta(x,y)$ (Eq. (7)) is the expectation of the product of $\phi$. Moreover, as $\phi=\text{erf},\sin,\cos$ and their corresponding NTKs are analytic, $f_{N}^{*}$ and $f_{\infty}^{*}$ have finitely many attractors, according to Theorem B in Ref. [41] (Sec. LABEL:S:sec:finiteness in the SM [26], which includes Ref. [42]). In contrast, certain trajectories for $\phi=\text{ReLU}$, whose NTK is unbounded, will head toward infinity.
@@
Figure 2: Change in dynamical system $f_{\infty}^{*}$ for $\phi=\text{erf}$ with respect to $c$, with $a=-1$, $b=1$, and $\sigma=1.0$. The trajectory $\{x_{n}\}_{n=0}^{T}$ of $T=10^{5}$ steps is computed with different $x_{0}$ and $c$ in given intervals, excluding $c=a,b$. (a) Changes in the acquired periods $\{x_{p}\}$ of period $p$, calculated by solving $(f_{\infty}^{*})^{p}(x_{p})=x_{p}$ (see Sec. LABEL:S:sec:LPn in the SM [26] for details). (b) Bifurcation diagram of the characteristic attractors calculated with $-10\leq x_{0}\leq 10$. The dotted lines indicate ${\cal D}$; the diagonal lines correspond to varying $c$. (c) Change in the basin of attraction where $f_{\infty}^{*}$ has multiple untrained attractors. (d) Change in the map $f_{\infty}^{*}$ in $-0.8\leq c\leq 0.8$. The circles indicate ${\cal D}$ and the red circle indicates a c-independent point $(a,b)=(-1,1)$. As $c$ approaches $a$ or $b$, the folding of $f_{\infty}^{*}$ around ${\cal D}$ becomes larger, making the characteristic attractors wider.